OpenStudy (muddygirl):

Help!! Solve this equation: 2(x + 6) = x + 17 A. x = 5 B. x = 11 C. x = 17 D. x = -1

OpenStudy (wendy.ivette11714):

2(x+6)=x+17 Step 1: Simplify both sides of the equation. 2(x+6)=x+17 Simplify: (2)(x)+(2)(6)=x+17(Distribute) 2x+12=x+17 Step 2: Subtract x from both sides. 2x+12−x=x+17−x x+12=17 Step 3: Subtract 12 from both sides. x+12−12=17−12 x=5

OpenStudy (muddygirl):

Thank you @Wendy.Ivette11714

OpenStudy (wendy.ivette11714):

Your Welcome :)